# Surgical-LVLM: Learning to Adapt Large Vision-Language Model for Grounded Visual Question Answering in Robotic Surgery

### v2

@@ IV-B Implementation Details
Surgical-LVLM 72.48 69.62
74.61 70.20
88.53 81.23
@@
TABLE II: Comparison experiments of our Surgical-LVLM against LVLMs without fine-tuning on EndoVis Conversations Dataset.
@@
BLIP-2 [35] 65.21 60.43
miniGPT-4 [36] 70.79 66.51
LLaVA-1.5-7B [23] 80.20 76.11
LLaVA-1.5-13B [23] 81.29 77.59
LLaVA-Med [37] 83.54 80.63
@@
TABLE III: Quantitative evaluation experiments on EndoVis-18 and EndoVis-17 datasets.
@@
VisualBERT [2] 0.6268 0.3329 0.7391 0.4005 0.3381 0.7073
VisualBERT R [3] 0.6301 0.339 0.7352 0.419 0.337 0.7137
MCAN [39] 0.6285 0.3338 0.7526 0.4137 0.2932 0.7029
VQA-DeiT [40] 0.6104 0.3156 0.7341 0.3797 0.2858 0.6909
MUTAN [39] 0.6283 0.3395 0.7639 0.4242 0.3482 0.7218
MFH [41] 0.6283 0.3254 0.7592 0.4103 0.35 0.7216
BlockTucker [42] 0.6201 0.3286 0.7653 0.4221 0.3515 0.7288
CAT-ViL DeiT [6] 0.6452 0.3321 0.7705 0.4491 0.3622 0.7322
GVLE-LViT [1] 0.6659 0.3614 0.7625 0.4576 0.2489 0.7275
@@
TABLE IV: Ablation study on different modified modules. ‘VP-LoRA’ denotes ‘Visual Perception LoRA’ and ‘MA’ denotes ‘Multimodal Alignment’.
@@
Surgical-LVLM 0.6878 0.3211 0.8303 0.4054 0.3386 0.7689
0.6947 0.3325 0.8322 0.4068 0.3412 0.7719
0.6878 0.3211 0.8336 0.4054 0.3386 0.7781
@@ IV-C Results
We first evaluate the effectiveness of our proposed methods on the EndoVis Conversation dataset in Table I. The Surgical-LVLM framework is evaluated based on whether VP-LoRA is introduced and whether it undergoes Instruction Fine-tuning. We can find that Instruction Fine-tuning is crucial for logical reasoning and comprehensive responses in surgical scenarios for the large language model. This is demonstrated by the huge gap in the GPT-4 Score. Therefore, the generalized large language models are not yet fully applicable in specialized scenarios such as surgery. Moreover, the introduction of VP-LoRA brings an effective enhancement to the model’s language response, proving its effectiveness in accurate response generation.
@@
Next, we compared the performance of our Surgical-LVLM framework against excellent LVLMs without Instruction Fine-tuning in surgical scenarios, which is shown in Table II. Significantly, Surgical-LVLM has demonstrated remarkable performance with the highest GPT-4 Score recorded at 90.68 on the EndoVis-18-VQLA dataset and 83.24 on the EndoVis-17-VQLA dataset, surpassing all competing models evaluated on the EndoVis Conversations Dataset. This notable margin of improvement further underscores the imperative for instruction tuning within LVLMs frameworks when applied to specific domains.
@@
Quantitative evaluation in Table III presents that our Surgical-LVLM outperforms other state-of-the-art models in traditional Surgical-VQLA tasks. On the EndoVis-18 dataset, our Surgical-LVLM model outperforms the others in terms of Accuracy and mIoU but has a lower F-Score. This indicates it is better at identifying the correct answers and detecting the images but might struggle slightly with balancing precision and recall. Its high mIoU suggests strong grounding capabilities. On the EndoVis-17 dataset, the performance of all models decreased due to suffering from domain migration. Despite the overall lower scores compared to EndoVis-18, the Surgical-LVLM still has the highest mIoU, confirming its robustness. However, other models, such as CAT-ViL DeiT and BlockTucker, have slightly better Accuracy and F-scores, hinting at potential areas of improvement for the Surgical-LVLM in terms of Accuracy and F-scores. Overall, the Surgical-LVLM model with Qwen large language model as well as multi-model alignment of CAT-ViL, which has excellent reasoning ability as well as strong inference-based grounding capability, as evidenced by mIoU metrics that are far superior to those of other state-of-the-art models.
@@
Furthermore, an ablation study is conducted to observe the effects of the designed modules in EndoVis-18-VQLA and EndoVis-17-VQLA datasets. Surgical-LVLM is categorized by the inclusion of VP-LoRA and Multimodal Alignment and the results are shown in Table IV. Incorporating VP-LoRA consistently enhances performance across both datasets and metrics. While MA shows improvement, its impact is focused on detection performance. Models combining both VP-LoRA and MA demonstrate the highest performance, indicating synergistic effects. Overall, these findings emphasize the importance of VP-LoRA in the propagation of global contextual information for accurate response generation, with complementary benefits when combined with MA for grounding.
IV-D Discussion and Limitations
@@
However, performance degradation across all models on the EndoVis-17 dataset underscores the challenge of maintaining robustness across different domains. While the Surgical-LVLM outperformed on various metrics, its slightly lower F-Score on the EndoVis-18 dataset suggests room for improvement in balancing precision and recall. In addition, further research could explore VP-LoRA’s optimal integration and potential limitations, although the its introduction enhances response accuracy. Finally, the deployment of LVLMs in real surgical scenarios is still limited by safety concerns due to the need to review the reliability in unpredictable surgical environments for potential safety risks.
V Conclusions
In this paper, we present Surgical-LVLM, a personalized vision-language model specifically designed for efficient grounded visual question answering in robotic surgery. The Surgical-LVLM leverages the specifically designed visual Perception LoRA to enable the model to understand complex visual-language surgical tasks. Additionally, we introduce a novel token interaction module to facilitate the critical language information emphasis. Our experimental results, conducted on multiple benchmarks, manifest the superiority of our proposed approach. However, deploying LVLMs in real-world surgical scenarios remains a long way off due to the need to thoroughly review the reliability and decision-making process of LVLMs in dynamic and unpredictable surgical environments to prevent potential risks. Therefore, future works shall highlight the need for ongoing research, ethical considerations, and robust safety mechanisms to realize the potential of foundation models in surgical applications.

### v3

@@ 4.2 Implementation Details
Surgical-LVLM $\times$ $\times$ 72.48 69.62
$\times$ 74.61 70.20
$\times$ 88.53 81.23
@@
Table 2: Comparison experiments of our Surgical-LVLM against LVLMs without fine-tuning on EndoVis Conversations Dataset.
@@
BLIP-2 (Li et al., 2023b) 65.21 60.43
miniGPT-4 (Zhu et al., 2023) 70.79 66.51
LLaVA-1.5-7B (Liu et al., 2023a) 80.20 76.11
LLaVA-1.5-13B (Liu et al., 2023a) 81.29 77.59
LLaVA-Med (Li et al., 2024a) 83.54 80.63
@@
Table 3: Quantitative evaluation experiments on EndoVis-18 and EndoVis-17 datasets.
@@
VisualBERT (Li et al., 2019) 0.6268 0.3329 0.7391 0.4005 0.3381 0.7073
VisualBERT R (Seenivasan et al., 2022a) 0.6301 0.339 0.7352 0.419 0.337 0.7137
MCAN (Ben-Younes et al., 2017) 0.6285 0.3338 0.7526 0.4137 0.2932 0.7029
VQA-DeiT (Touvron et al., 2021) 0.6104 0.3156 0.7341 0.3797 0.2858 0.6909
MUTAN (Ben-Younes et al., 2017) 0.6283 0.3395 0.7639 0.4242 0.3482 0.7218
MFH (Yu et al., 2018) 0.6283 0.3254 0.7592 0.4103 0.35 0.7216
BlockTucker (Ben-Younes et al., 2019) 0.6201 0.3286 0.7653 0.4221 0.3515 0.7288
CAT-ViL DeiT (Bai et al., 2023b) 0.6452 0.3321 0.7705 0.4491 0.3622 0.7322
GVLE-LViT (Bai et al., 2023d) 0.6659 0.3614 0.7625 0.4576 0.2489 0.7275
@@
Table 4: Ablation study on different modified modules. ‘VP-LoRA’ denotes ‘Visual Perception LoRA’ and ‘MA’ denotes ‘Multimodal Alignment’.
@@
Surgical-LVLM $\times$ $\times$ 0.6878 0.3211 0.8303 0.4054 0.3386 0.7689
$\times$ 0.6947 0.3325 0.8322 0.4068 0.3412 0.7719
$\times$ 0.6878 0.3211 0.8336 0.4054 0.3386 0.7781
@@ 4.3 Results
We first evaluate the effectiveness of our proposed methods on the EndoVis Conversation dataset in Table 1. The Surgical-LVLM framework is evaluated based on whether VP-LoRA is introduced and whether it undergoes Instruction Fine-tuning. We can find that Instruction Fine-tuning is crucial for logical reasoning and comprehensive responses in surgical scenarios for the large language model. This is demonstrated by the huge gap in the GPT-4 Score. Therefore, the generalized large language models are not yet fully applicable in specialized scenarios such as surgery. Moreover, the introduction of VP-LoRA brings an effective enhancement to the model’s language response, proving its effectiveness in accurate response generation.
@@
Next, we compared the performance of our Surgical-LVLM framework against excellent LVLMs without Instruction Fine-tuning in surgical scenarios, which is shown in Table 2. Significantly, Surgical-LVLM has demonstrated remarkable performance with the highest GPT-4 Score recorded at 90.68 on the EndoVis-18-VQLA dataset and 83.24 on the EndoVis-17-VQLA dataset, surpassing all competing models evaluated on the EndoVis Conversations Dataset. This notable margin of improvement further underscores the imperative for instruction tuning within LVLMs frameworks when applied to specific domains.
@@
Quantitative evaluation in Table 3 presents that our Surgical-LVLM outperforms other state-of-the-art models in traditional Surgical-VQLA tasks. On the EndoVis-18 dataset, our Surgical-LVLM model outperforms the others in terms of Accuracy and mIoU but has a lower F-Score. This indicates it is better at identifying the correct answers and detecting the images but might struggle slightly with balancing precision and recall. Its high mIoU suggests strong grounding capabilities. On the EndoVis-17 dataset, the performance of all models decreased due to suffering from domain migration. Despite the overall lower scores compared to EndoVis-18, the Surgical-LVLM still has the highest mIoU, confirming its robustness. However, other models, such as CAT-ViL DeiT and BlockTucker, have slightly better Accuracy and F-scores, hinting at potential areas of improvement for the Surgical-LVLM in terms of Accuracy and F-scores. Overall, the Surgical-LVLM model with Qwen large language model as well as multi-model alignment of CAT-ViL, which has excellent reasoning ability as well as strong inference-based grounding capability, as evidenced by mIoU metrics that are far superior to those of other state-of-the-art models.
@@
Furthermore, an ablation study is conducted to observe the effects of the designed modules in EndoVis-18-VQLA and EndoVis-17-VQLA datasets. Surgical-LVLM is categorized by the inclusion of VP-LoRA and Multimodal Alignment and the results are shown in Table 4. Incorporating VP-LoRA consistently enhances performance across both datasets and metrics. While MA shows improvement, its impact is focused on detection performance. Models combining both VP-LoRA and MA demonstrate the highest performance, indicating synergistic effects. Overall, these findings emphasize the importance of VP-LoRA in the propagation of global contextual information for accurate response generation, with complementary benefits when combined with MA for grounding.
4.4 Discussion and Limitations
@@
However, performance degradation across all models on the EndoVis-17 dataset underscores the challenge of maintaining robustness across different domains. While the Surgical-LVLM outperformed on various metrics, its slightly lower F-Score on the EndoVis-18 dataset suggests room for improvement in balancing precision and recall. In addition, further research could explore VP-LoRA’s optimal integration and potential limitations, although its introduction enhances response accuracy. Finally, the deployment of LVLMs in real surgical scenarios is still limited by safety concerns necessitating a rigorous evaluation of their reliability in complex surgical environments to mitigate potential risks.
5 Conclusions
In this paper, we present Surgical-LVLM, a personalized vision-language model specifically designed for efficient grounded visual question answering in robotic surgery. The Surgical-LVLM uses the specifically designed VP-LoRA to enable the model to understand complex visual-language surgical tasks. Additionally, we introduce a novel token interaction module to facilitate the critical language information emphasis. Our experimental results, conducted on multiple benchmarks, manifest the superiority of our proposed approach. However, deploying LVLMs in real-world surgical scenarios remains a long way off due to the need to thoroughly review the reliability and decision-making process of LVLMs in dynamic and unpredictable surgical environments. Therefore, future works shall highlight the need for ethical considerations, and robust safety mechanisms to realize the potential of foundation models in surgical applications.
Acknowledgments
This work was supported by Hong Kong RGC CRF C4026-21GF, GRF 14203323, GRF 14216022, GRF 14211420, NSFC/RGC Joint Research Scheme N_CUHK420/22, Shenzhen-Hong Kong-Macau Technology Research Programme (Type C) STIC Grant 202108233000303.